OpenStudy (anonymous):

Which answer shows this equation in standard form? 7 – 3(x – y) = –5x + 2

OpenStudy (igreen):

Standard form is \(Ax + By = C\). First distribute \(-3\) into the parenthesis. \(7 - 3(x - y) = -5x + 2\) \(7 - 3x + 3y = -5x + 2\) Can you do the rest? @OGangsta

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

7-(-2)+3y=2???

OpenStudy (igreen):

No, first you what to subtract 7 to both sides..because you want to keep the regular whole numbers without variables to the right of the equal sign.

OpenStudy (anonymous):

oh i subtracted -5x

OpenStudy (igreen):

You're actually supposed to add 5x..and I don't see any term with 'x' in \(7-(-2)+3y=2\).

OpenStudy (anonymous):

thats what i have

OpenStudy (igreen):

I know..and I don't see any 'x' term in it..

OpenStudy (igreen):

\(7 - 3x + 3y = -5x + 2\) Okay, now we subtract \(7\) to both sides: \( \cancel{7 - 7} -3x + 3y = -5x + 2 - 7\) \(-3x + 3y = -5x - 5\) Now, add \(5x\) to both sides.

OpenStudy (anonymous):

8x+3y=5

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

-8x+3y=-5

OpenStudy (igreen):

No, that's close \(3y\) and \(-5\) are correct..but \(5x + -3x\) isn't \(8x\) or \(-8x\), it's the same thing as \(5x - 3x\). @OGangsta

OpenStudy (anonymous):

so 2x

OpenStudy (igreen):

Yep, so what's your final answer going to be?

OpenStudy (anonymous):

2x+3y=-5

OpenStudy (igreen):

Yep!

OpenStudy (anonymous):

thank you